# A Stochastic–Geometrical Framework for Object Pose Estimation based on Mixture Models Avoiding the Correspondence Problem

### v1

@@ 3.2 Visualization of the Objective Function
In order to achieve a reasonable visual comparison, we utilize the described object with the same rigid body transformation. The center of the visual comparison is the plot of the objective function, which is generated for all three systems under varying circumstances. In order to achieve a direct impression of the objective function, we utilize a pose transformation which consists of one rotation angle $\varphi$ and one translational parameter $w$ and which, in consequence, can be plotted as a 2D image in the $(\varphi,w)$–plane. With these assumptions, typical characteristics of interest in this paper can be effectively illustrated. The rigid body transformation depending on the parameters $\vec{\beta}:=(\varphi,w)$ is defined by
$\vec{p}_{1}(\vec{z};(\varphi,w)):=\left(\begin{array}[]{c}\cos(\varphi)\cdot(z_{1}+w)+\sin(\varphi)\cdot(z_{2}+w)\\ -\sin(\varphi)\cdot(z_{1}+w)+\cos(\varphi)\cdot(z_{2}+w)\\ z_{3}+w\end{array}\right)\;.$
The observational function $\vec{p}_{2}$ (recall, $\vec{p}(\vec{z};\vec{\beta},\vec{\gamma}):=\vec{p}_{2}(\vec{p}_{1}(\vec{z};\vec{\beta});\vec{\gamma})$ from subsection 2.1.2) for the first two systems is the same (orthogonal parallel beam camera projection) as described in the camera application subsection 2.3.2. For the third system (lateration) we utilize the observation function as described in lateration example application subsection 2.3.3.
@@ 3.3 Accuracy Evaluation
For demonstration purposes, we will randomly sample the true values $(\vec{\varphi},\vec{w})$ for each simulation run from a uniform distribution on the intervals $\varphi_{1},\varphi_{2}\in[-\pi/36,\pi/36]$ and $\varphi_{3}\in[-\pi/3-\pi/36,-\pi/3+\pi/36]$, as well as $w_{1},w_{2},w_{3}\in[35,45]$ ($\pi/36$ corresponds to $5^{\circ}$, which means in this simulation a ±$\pm 5^{\circ}$ variation of each true angular value is illustrated, which is regarded as a significant range). These intervals correspond to a similar setting as in subsection 3.2 with additional variations of the presented objective functions in figure 3, in order to provide a continuation to the previous results.
The number of samples in the Monte–Carlo integration is selected for system 2 and 3 to be constant by $R=1000$ and chosen by a normally distributed latin hypercube sampling by Matlab’s lhsnorm–function.
For the realization of the optimization in each simulation run a standard routine is selected with the Nelder–Mead–Algorithm of Matlab’s fminsearch–function. The starting values of this local optimization are selected also randomly by varying the true values $\vec{\varphi}+\Delta\vec{\varphi}$ and $\vec{w}+\Delta\vec{w}$ with an additional uniform sampling from the intervals $\Delta\varphi_{1},\Delta\varphi_{2},\Delta\varphi_{3}\in[-\pi/72,\pi/72]$ (i.e. $[-2.5^{\circ},2.5^{\circ}]$) and $\Delta w_{1},\Delta w_{2},\Delta w_{3}\in[-5,5]$. The overall random selection of the true values for each simulation run as well as the random selection of starting values for the optimization has the goal to demonstrate the effectiveness and overall accuracy of the algorithm for the three different systems under varying situations.
@@
system 1: $I$ 0.00892 (0.511) 0.00923 (0.529) 0.00313 (0.179) 0.34425 0.28588 1.27141
$II$ 0.01463 (0.838) 0.02061 (1.181) 0.01406 (0.805) 1.02502 1.38896 2.78294
$III$ 0.00774 (0.444) 0.01222 (0.700) 0.00359 (0.206) 0.32492 0.30322 1.35315
$IV$ 0.02564 (1.469) 0.02561 (1.468) 0.02311 (1.324) 2.59090 2.69530 2.53232
system 2: $I$ 0.00007 (0.004) 0.00007 (0.004) 0.00004 (0.002) 0.04534 0.04651 0.04394
$II$ 0.01264 (0.724) 0.02158 (1.237) 0.01479 (0.847) 1.07286 1.45570 2.92436
$III$ 0.00006 (0.003) 0.00006 (0.004) 0.00004 (0.002) 0.04731 0.04410 0.04568
$IV$ 0.00982 (0.563) 0.01002 (0.574) 0.00532 (0.305) 0.56681 0.56640 1.30087
system 3: $I$ 0.02377 (1.362) 0.02161 (1.238) 0.01994 (1.142) 1.64819 1.71159 2.76714
$II$ 0.20114 (11.525) 0.19631 (11.248) 0.18541 (10.623) 11.18420 16.77133 13.31179
$III$ 0.01144 (0.656) 0.01975 (1.132) 0.01490 (0.854) 1.14116 1.36369 2.42605
$IV$ 0.03018 (1.729) 0.02815 (1.613) 0.02191 (1.255) 2.02768 1.97756 3.28037
Table 2: Presentation of the statistical evaluation of the full $6$D pose estimation results for system 1, 2 and 3, as well as for each simulation scenario $I$$IV$. The residual errors are presented as RMS–values of all $6$D parameters with the angular units $\varphi_{i}$ in [rad] (and [deg] in brackets) and $w_{i}$ in the length unit.
@@
Second, system 2 outperforms system 3 in essentially all simulation scenarios, which means that under the given parameter set the two orthogonal camera system is in general better than the orthogonal trilateration system. This is in accordance to the observations in subsection 3.2, which shows a less distinct maximum for system 3 than for system 2 in the objective function. It has to be noted that this does not mean, that the trilateration system could not be better in practice, since the parameter of distance measurement accuracy $\sigma_{\varepsilon}$ could be much lower for lateration systems (based on laser interference) relative to camera systems with given pixel resolutions. It is pointed to the fact, that within this unifying modeling framework it is possible to determined how much the distance measurement accuracy in the lateration system must be better to outperform the camera setup.
Third, scenario $II$ shows that for low–resolution measurements all systems 1–3 are affected strongly, even to the point of uselessness of the estimation. If the inherent resolution of the measurement equipment is too low for the $6$D pose estimation, a different placement of the measurement devices or an increased number of measurement devices could be investigated within this framework.
@@ 4 Discussion
From a practitioner’s perspective, defining the necessary density functions in this new framework might be challenging since each of them should be backed up by independent observations, e.g. the feature extraction error density or realistic feature position densities in object space. However, this effort has its own merit, since this new type of blurry information is transparently incorporated into the estimation process applying this framework (which benefit was presented in the numerical results section). If these independent pieces of information are not available, reasonable assumptions must be defined. It is also important to notice, that finding advantageous geometric parameter sets $\vec{\gamma}$ in the observation model might be challenging, especially, if one wants to define which parameters are set to fixed values and which are optimized in a calibration routine of a multi–view (and possibly multi–technological) system.
@@
In this paper we propose a model driven first principles approach by defining random variables and show how they are transformed in an observation scenario to estimate an object pose. We acknowledge that this in the contrast to the algorithmically driven current main body of research based on (deep learning) artificial neural networks architectures. We want to stress that these papers are very successful and mainly rely on an extensive training for specific camera setups and exemplary video stream data which is shared in the computer vision community (e.g. Kendall et al.[16]) and Rad et al.[19]). Compared to this, the goal of this paper is to allow to design and compare new observation scenarios and devices. In consequence, we do not think that it is reasonable to compare the general argumentation framework resulting in a general algorithmic family (with plausibility simulations in the results section) of this presentation to specifically highly trained algorithmical approaches on example video stream data sets which are generated by given camera devices. On the one hand, we do encourage future work in developing and adapting the presented algorithmic family further to specific applications so that comparison to existing algorithmical approaches can be reasonably performed. On the other hand, we are optimistic that combining the probabilistic view of this modeling framework with algorithm driven solutions can lead to improvements of pose estimation in the future.

### v4

@@ 3.2 Visualization of the Objective Function
In order to achieve a reasonable visual comparison, we utilize the described object with the same rigid body transformation. The center of the visual comparison is the plot of the objective function, which is generated for all three systems under varying circumstances. In order to achieve a direct impression of the objective function, we utilize for all systems the diagonally shifted and around $z_{3}$–axis rotated transformation $\vec{p}_{1}(\vec{z};(\varphi,w))$ as presented in subsection 2.3.2. For the first two systems the same (orthogonal parallel beam camera projection) $\vec{p}_{2}$ is utilized as described in the camera application subsection 2.3.2. For the third system (lateration) we utilize the observation function $p_{2}$ as described in lateration example application subsection 2.3.3.
@@ 3.3 Accuracy Evaluation
For demonstration purposes, we will randomly sample the true values $(\vec{\varphi},\vec{w})$ for each simulation run from a uniform distribution on the intervals $\varphi_{1},\varphi_{2}\in[-\pi/36,\pi/36]$ and $\varphi_{3}\in[-\pi/3-\pi/36,-\pi/3+\pi/36]$, as well as $w_{1},w_{2},w_{3}\in[35,45]$ ($\pi/36$ corresponds to $5^{\circ}$, which means in this simulation a ±$\pm 5^{\circ}$ variation of each true angular value is illustrated, which is regarded as a significant range relative to the expected accuracy). These intervals correspond to a similar setting as in subsection 3.2 with additional variations of the presented objective functions in figure 5, in order to provide a continuation to the previous results.
For the realization of the optimization in each simulation run a standard routine is selected with the Nelder–Mead–Algorithm of Matlab’s fminsearch–function. The starting values of this local optimization are selected also randomly by varying the true values $\vec{\varphi}+\Delta\vec{\varphi}$ and $\vec{w}+\Delta\vec{w}$ with an additional uniform sampling from the intervals $\Delta\varphi_{1},\Delta\varphi_{2},\Delta\varphi_{3}\in[-\pi/72,\pi/72]$ (i.e. $[-2.5^{\circ},2.5^{\circ}]$) and $\Delta w_{1},\Delta w_{2},\Delta w_{3}\in[-5,5]$. The random selection of the true values for each simulation run as well as the random selection of starting values for the optimization has the goal to demonstrate the effectiveness and overall accuracy of the algorithm for the three different systems under varying situations.
Regarding the computational effort, the number of samples in the Monte–Carlo integration is selected for system 2 and 3 to be constant with $R=1000$ and sampled by a normally distributed latin hypercube sampling by Matlab’s lhsnorm–function. In total, for the implementation a standard tooling is utilized with only a moderate focus on a computational runtime. In order to judge the computational efficiency, the most runtime is due to interpolation of the observations $g_{l}$. For system 2 a 2D bilinear interpolation routine is utilized for evaluating the two 2D images $g_{1}$ and $g_{2}$, and for system 3 a 1D linear interpolation is utilized for evaluating the three 1D signals $g_{1},g_{2},g_{3}$. The runtime for one evaluation of the objective function (recall, with $R=1000$ interpolations per feature and observation) for system 2 is on average $10.8$ milliseconds and for system 3 on average $10.2$ milliseconds on a standard working laptop with computation on a $1.8$ Ghz single core. For one $6$D pose estimation with the Nelder–Mead–Algorithm in the most demanding scenario $IV$ approximately $300-1200$ objective function evaluations are necessary (strongly depending on each simulation run), leading to computational times between $3-13$ seconds per pose estimation. It must be noticed that in this evaluation it is not the focus to present the most efficient implementation of the proposed optimization problem but to show the typical accuracy behavior in a simulation study depending on scenarios $I-IV$. In consequence, we expect several factors of acceleration in computational runtime with high efficiency programming, possibly allowing real–time tracking applications.
@@
system 1: $I$ 0.51 0.53 0.18 0.344 0.286 1.271
$II$ 0.84 1.18 0.81 1.025 1.389 2.783
$III$ 0.44 0.70 0.21 0.325 0.303 1.353
$IV$ 1.47 1.47 1.32 2.591 2.695 2.532
system 2: $I$ 0.004 0.004 0.002 0.045 0.047 0.044
$II$ 0.72 1.24 0.85 1.073 1.456 2.924
$III$ 0.003 0.004 0.002 0.047 0.044 0.047
$IV$ 0.56 0.57 0.31 0.567 0.566 1.301
system 3: $I$ 1.36 1.24 1.14 1.648 1.712 2.767
$II$ 11.53 11.25 10.62 11.184 16.771 13.312
$III$ 0.67 1.13 0.85 1.141 1.364 2.426
$IV$ 1.73 1.61 1.26 2.028 1.978 3.280
Table 2: Presentation of the statistical evaluation of the full $6$D pose estimation results for system 1, 2 and 3, as well as for each simulation scenario $I$$IV$. The residual errors are presented as RMS–values of all $6$D parameters with the angular units $\varphi_{i}$ in [deg] and $w_{i}$ in the length unit.
@@
Second, system 2 outperforms system 3 in essentially all simulation scenarios, which means that under the given parameter set the two orthogonal camera system is in general better than the orthogonal trilateration system. This is in accordance to the observations in subsection 3.2, which shows a less distinct maximum for system 3 than for system 2 in the objective function. It has to be noted that this does not mean that the trilateration system could not be better in practice, since the parameter of distance measurement accuracy $\sigma_{\varepsilon}$ could be much lower for lateration systems (based on laser interference) relative to camera systems with given pixel resolutions. It is pointed to the fact, that within this unifying modeling framework it is possible to determine how much the distance measurement accuracy in the lateration system must be better to outperform the camera setup.
Third, scenario $II$ shows that for low–resolution measurements all systems 1–3 are affected strongly, even to the point of uselessness of the estimation. If the inherent resolution of the measurement equipment is too low for the $6$D pose estimation, a different placement of the measurement devices or an increased number of measurement devices could be investigated within this framework. A general tipping point for low–resolution measurements is that the individual point features cannot be separated anymore in the observation. This depends, of course, on the physical distance of the features and their arrangement on the object. Such a problematic situation is demonstrated in figure 4 where the individual feature observations are not possible anymore in scenario $II$ for system 3, leading to the worst pose estimation results in this quantitative analysis.
@@ 4 Discussion
From a practitioner’s perspective, defining the necessary density functions in this new framework might be challenging since each of them should be backed up by independent observations, e.g. the feature extraction error density or realistic feature position densities in object space. However, this effort has its own merit, since this new type of blurry information is transparently incorporated into the estimation process applying this framework (which benefit was presented in the numerical results section). If these independent pieces of information are not available, reasonable assumptions must be defined such as working with specification sheets of system components (e.g. point spread functions of a camera lens) or typically expected uncertainties in comparable observation scenarios. In such situations, it is in general difficult to assess the influence of modeling errors since it is highly application dependent. Since this approach is focusing on a stochastic best fit, minor modeling errors in the densities are expected to have only minor influence on the pose estimation. It is also important to notice, that finding advantageous geometric parameter sets $\vec{\gamma}$ in the observation model might be challenging, especially, if one wants to define which parameters are set to fixed values and which are optimized in a calibration routine of a multi–view (and possibly multi–technological) system.
@@
In this paper we propose a model driven first principles approach by defining random variables and show how they are transformed in an observation scenario to estimate an object pose. We acknowledge that this in the contrast to the algorithmically driven current main body of research, which is based on (deep learning) artificial neural networks architectures. We want to stress that these papers are very successful and mainly rely on an extensive training for specific camera setups and exemplary video stream data which is shared in the computer vision community (e.g. Kendall et al.[16]) and Rad et al.[19]). Compared to this, the goal of this paper is to allow to design and compare new observation scenarios and devices. In consequence, we do not think that it is reasonable to compare the general argumentation framework resulting in a general algorithmic family (with plausibility simulations in the results section) of this presentation to specifically highly trained algorithmical approaches on example video stream data sets which are generated by given camera devices. On the one hand, we do encourage future work in developing and adapting the presented algorithmic family further to specific applications so that comparison to existing algorithmical approaches can be reasonably performed. On the other hand, we are optimistic that combining the probabilistic view of this modeling framework with algorithm driven solutions can lead to improvements of pose estimation in the future.